OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (anonymous):

okay let's focus on the second equation first. x+3y=9 let's get x by it self x=-3y+9 now. place this in for y in the first equation and solve for y (-3y+9)-2y=14 simplify -5y+9=14 -9 from both sides -5y=5 divide by -5 y=-1 plug this in ether equation so I'm going to place it in the second one x+3(-1)=9 simplify x+ -3=9 +3 both side x=12 that means the answer is x=12 y=-1 (12,-1) Hope this helps!

OpenStudy (anonymous):

thank you so much :3 can u help me with 2 more ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

and thx :3

OpenStudy (anonymous):

okay here is this one. start with the first equation. get y by itself y=2x+1 plug that in to the second question -4x+(2x+1)=-1 simplify and solve for x 2x+1=-1 2x=-2 x=-1 plug this in to any equation i'm going to go with the second one -4x+y=-1 -4(-1)+y=-1 4+y=-1 y=5....hmmm

OpenStudy (anonymous):

oh ok thank you so much :3

OpenStudy (anonymous):

it's (1,3)

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

u think so ^-^ thank you so much for your help :)

OpenStudy (anonymous):

haha your welcome :) Good luck!

OpenStudy (anonymous):

thank ^-^